OpenStudy (anonymous):

How would you solve x^4=2x^2 - 1?

OpenStudy (anonymous):

still not understanding..

OpenStudy (campbell_st):

rewrite it as \[x^4 - 2x^2 + 1 = 0\] let \[u = x^2\] solve u^2 -2u+ 1 = 0 which is (u - 1)^2 = 0 then u = 1 is a solution but since u = x^2 then x^2 = 1 or \[x = \pm\]

OpenStudy (campbell_st):

oops \[x = \pm 1\]

OpenStudy (anonymous):

where'd u come from is what I'm curious of?

OpenStudy (anonymous):

IT MAKES SENSE NOW!!! THANKYOU